Project

General

Profile

20160503 » History » Version 12

Boyd, Steven, 03 May 2016 16:21

1 1 Boyd, Steven
h1. Ops/Comm Scheduling meeting
2
3
Location : CR11 in R3 @ 15:30 pm on Tuesday 3rd
4
5
Phone access :  ReadyTalk number 
6
US 866-740-1260
7
UK 0800-496-0576
8
Swiss 0800-70-53-51
9
Italy 800182592
10
Access Code: 2155402
11
12
h1. Agenda
13
14
* *Actions from last meeting*
15
** Generate appropriate table to store absorber pressure and material information  (Janusz) : (almost done)
16
** DS readback current requires calibration (DL) : stands until close to powerup
17
** Issues arising from last data taking run need to be identified and actioned (Boyd) : Stands
18
** Move the tracker CAENbus module higher up the chain (Craig)
19
** Get an estimate for how long it takes to wire in the chiller in (Andy)
20
21
* DC report
22 5 Boyd, Steven
* Work plan for the next week : http://micewww.pp.rl.ac.uk/attachments/6180/2016-05-03-MICE-Hall-Plan.xlsx
23 10 Boyd, Steven
* QP/QD work : http://micewww.pp.rl.ac.uk/attachments/6181/MICE_Electrical_Work_List_25APR16.docx
24 1 Boyd, Steven
* Water system work plan
25
* Bellows installation and FC cooldown
26
* Tracker hall probe installation
27
* AOB
28 4 Karadzhov, Yordan
** Prescaled trigger discussion (Online group). If you are interested, please indicate your availability in the doodle:
29 11 Karadzhov, Yordan
http://doodle.com/poll/wnkk4aevq938keb8
30 1 Boyd, Steven
31
* DONM : 10/3/2016
32 12 Boyd, Steven
h1. Minutes
33
34
* Present : Yordan, Paul H, Paul S, Mike C, Vicky B, Josef F, Coli W, Andy N, Craig M, Steve B, Pierrick H
35
36
* *Actions from last meeting*
37
** Generate appropriate table t 
38
39
* DC report
40
** Focus coil won't go back in since the bellows have been switched around and now the screw hole tolerances no longer match.
41
** When John gets back from holiday the FC will be pulled out and the bellows swapped
42
** Daresbury work went well. An ongoing work plan is shown below
43
** US team came out with a program of work, some of which can be done but because the racks are not fully commissioned the
44
full program of work can't be completed.
45
** This week HPC have started to work on the compressor for the DS. They are concerned about the temperature of the air end - it's
46
too hot to touch. They plan to replace this in August. Colin:Can we just buy a new one rather than replace in August? Mike: What we
47
get is effectively new but we should put an order in. Colin:we should do a comprehensive job on fixing the compressor so need to
48
understand this issue. Action Item: Vicky will investigate the state of the refurbished air end to see what we get for our money.
49
** Daresbury will allow the US to run into a short this week.
50
** Alan B : the 20 psi main relief valve for SSU needs replacing. Alan has a replacement but it needs installation. This can happen
51
when John is back next week.
52
* Work plan for the next week : http://micewww.pp.rl.ac.uk/attachments/6180/2016-05-03-MICE-Hall-Plan.xlsx
53
** Henry: Water plan goes into period when the DSA is closed. If we turn the water back on we may be in a situation where the
54
water pipes to the conventional magnets could blow. If the DSA is closed we can't get to it. Colin: Good point. Henry: We also
55
need the full ISIS running schedule spreadsheet. Colin:I've got something - I'll send it to you. Action Item: Colin will put the
56
ISIS dates into his schedules including run up.
57
* QP/QD work 
58
* Water system work plan : Colin:There is no day-by-day plan, but the plan for what is being done is understood. The plan is not going
59
to plan due to delays in getting an external contractor. We'll put a project plan up here.
60
* Bellows installation and FC cooldown : Colin:We can plan to pump the FC down and get it as dry as it can. Is there any reason
61
to start cooling once it's pumped down. Josef: we can do this. Colin:once we're happy then we can start the coolers. 
62
* Tracker hall probe installation: Alan: Will start tomorrow - we have a plan which will get the losses down to an acceptable level.
63
* AOB
64
** Prescaled trigger discussion: Yordan is organising a phone meeting to develop the prescaled trigger. For those interested there
65
is a Doodle poll for this meeting.
66
** Henry: Still looking for people to train as BLOCs. BLOC training will happen in June. An invitation will be given at the VC.
67
** Henry: Water detection in the trench? Pierrick: we have water detection in the trench but it's in the gutter and so shows water
68
detection all the time. We need to come up with a proper plan. Action Item: Henry, Craig, Paul and Pierrick to look into optimising
69
this for once and for all.
70
** Henry: Can Paul lower the target frame and check that the laser light levels haven't changed. Yes.